A can do a work in 32 days and B in 40 days. they work together for 10 days and then B left. In how many days A will finish the remaining work??
Answer by mananth(16949)

A can do a work in 32 days and B in 40 days. they work together for 10 days and then B left. In how many days A will finish the remaining work?
A does 1/32 of the job in 1 day
B does 1/40 of the job in 1 day
..
Together they do 1/32 +1/40 in 1 day
= 72/1280 of the job in 1 day.
In 10 days they do 72/1280 * 10
=72/128.
..
balance work will be 1- 72/128
128-72 / 128
=56 /128
=7/16
1/32 of the job A does in 1 day
7/16 of the job he will do 7/16 / 1/32
=7/16 *32
= 14 days
A will take 14 more days to complete the job.
...